OpenStudy (anonymous):

Can someone check my answer? Did I do the correct process? (x+2) (x-2) = 32 FOIL x^2 -2x +2x -4 =32 x^2 - 4 = 32 +4 x^2 = 36 x= 6

OpenStudy (anonymous):

Yep, you did everything fine :)

OpenStudy (anonymous):

Ja/Yes.

OpenStudy (foolaroundmath):

x can also be -6 ;) x^2 = 36 => (x^2 - 6^2) = 0 => (x+6)(x-6) = 0 so , x=6 or x=-6

OpenStudy (anonymous):

^good point.

OpenStudy (anonymous):

\(\pm 6\)

OpenStudy (anonymous):

±6